OpenStudy (anonymous):

(a) An angle measures 34 . What is the measure of its supplement? (b) An angle measures 35 . What is the measure of its complement?

OpenStudy (anonymous):

a) 180-34=146 b) 90-35= 55

OpenStudy (anonymous):

x+(x-34)=180 2x-34=180 2x=180+34 2x=214 x=214/2 x=107 answer for the smaller angle proof 107+(107-34)=180 107+73=180 180=180

OpenStudy (anonymous):

that is A